OpenStudy (anonymous):

How many solutions does this system of equations have? x+y= -4 x -y= 0

OpenStudy (anonymous):

How many solutions does this system of equations have? x+y= -4=x+y+4=0 y=-x-4 y=-x-4

OpenStudy (anonymous):

infinite.correct?

OpenStudy (anonymous):

??